# Yibin Wang ${}^{\text{\textdagger}1}$ &Haizhou Shi ${}^{*\text{\textdagger}1}$ &Ligong Han 12

### v3

@@ Abstract
Large Language Models (LLMs) often suffer from overconfidence during inference, particularly when adapted to downstream domain-specific tasks with limited data. Previous work addresses this issue by employing approximate Bayesian estimation after the LLMs are trained, enabling them to quantify uncertainty. However, such post-training approaches’ performance is severely limited by the parameters learned during training. In this paper, we go beyond post-training Bayesianization and propose Bayesian Low-Rank Adaptation by Backpropagation (BLoB), an algorithm that continuously and jointly adjusts both the mean and covariance of LLM parameters throughout the whole fine-tuning process. Our empirical results verify the effectiveness of BLoB in terms of generalization and uncertainty estimation, when evaluated on both in-distribution and out-of-distribution data.
@@ 1 Introduction
Despite the recent advancements in Large Language Models (LLMs) [9, 99, 98, 65, 15, 4, 88, 89, 76, 12, 1, 70], addressing the challenges of reliability and responsibility remains imperative [42]. LLMs often produce overconfident responses detached from factual grounding, posing potential harm to users [3, 100, 44, 43, 84, 50, 7, 110, 104, 112, 33, 67, 107, 45]. Therefore, accurately estimating response confidence (or uncertainty) is crucial to preemptively intervene before harm occurs. Current research predominantly focuses on eliciting the internal capability of uncertainty estimation of LLMs. For example, studies suggest that verbalized uncertainty yields better-calibrated results compared to conditional probability [84, 45].
While effective, the aforementioned methods do not offer a universal solution for expressing LLM uncertainty across all scenarios, especially when adapted to domain-specific corpora, human preferences, or downstream tasks [44]. Even a well-calibrated LLM may struggle to estimate uncertainty during fine-tuning due to catastrophic forgetting of general knowledge [82]. Moreover, when applied to limited-scale downstream tasks, excessively over-parameterized LLMs can rapidly overfit, leading to overconfidence. Thus, enabling accurate uncertainty estimation of LLMs is vital for their reliable and responsible deployment.
Bayesian methods emerge as a natural solution for learning uncertainty estimation abilities among their counterparts [85, 11, 94, 29, 46, 51, 61, 58, 95, 20, 103]. These methods model predictive uncertainty $P({\bm{y}}|{\bm{x}},{\mathcal{D}})$ by marginalizing the posterior parameter distribution $P({\bm{\theta}}|{\mathcal{D}})$ after observing the dataset ${\mathcal{D}}$:
@@ 2.2 Variational Bayesian Networks (VBNs)
$\min_{{\bm{\theta}}}{\mathcal{F}}({\mathcal{D}},{\bm{\theta}})\triangleq-\mathbb{E}_{q({\bm{W}}|{\bm{\theta}})}[\log P({\mathcal{D}}|{\bm{W}})]+\operatorname{KL}[q({\bm{W}}|{\bm{\theta}})\parallel P({\bm{W}})].$ (3)
@@
${\mathcal{F}}({\mathcal{D}},{\bm{\theta}})$ $\approx-\tfrac{1}{n}\sum_{i=1}^{n}\log P({\mathcal{D}}|{\bm{W}}_{i})+\tfrac{1}{n}\sum_{i=1}^{n}[\log q({\bm{W}}_{i}|{\bm{\theta}})-\log P({\bm{W}}_{i})],$ (4)
where ${\bm{W}}_{i}={\bm{\mu}}+\log(1+\exp({\bm{\rho}}))\odot{\bm{\epsilon}}_{i}$ is the $i$-th sample of the weights yielded by parameterization and ${\bm{\epsilon}}_{i}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$. In BBB, the authors assume the prior distribution $P({\bm{W}})=\pi{\mathcal{N}}({\bm{0}},{\bm{\sigma}}_{1}^{2})+(1-\pi){\mathcal{N}}({\bm{0}},{\bm{\sigma}}_{2}^{2})$ to be a mixture of Gaussians. Consequently, they optimize the second term based on weight sampling. In different scenarios, a simpler form of the prior, which allows for a closed-form solution, can also be considered. Although our proposed method is largely based on the existing framework of BBB, trivially combining BBB with LoRA does not yield satisfactory results. It is important to note that our specific designs are necessary to encourage the fast convergence of the variational distribution, which will be introduced later in Sec. 3.
@@ 3.1 Low-Rank Variational Approximate Posterior Distribution: LoRA Bayesianization
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}]\cdot[\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\top}].$ (8)
Theorem 3.1 shows that our asymmetric LoRA Bayesianization is equivalent to using a Gaussian variational distribution for the full weight ${\bm{W}}$ (i.e., Eqn. 6), with a flexible covariance matrix (i.e., Eqn. 8), to approximate the postetior distribution of the full weight ${\bm{W}}$.
@@ 3.2 Low-Rank Prior Distribution
Theorem 3.2 shows that with a proper $\widetilde{{\bm{R}}}$, one can compute the KL divergence for the high-dimensional full weight $\operatorname{vec}({\bm{W}})$ simply by computing the KL divergence for ${\bm{A}}$, which is much lower-dimension, more parameter-efficient, more memory-efficient, and faster.
@@
$\min\operatorname{KL}[q({\bm{A}}|{\bm{\theta}}=\{{\bm{M}},{\bm{\Omega}}\})\|P({\bm{A}})]$ $=\tfrac{1}{2\sigma_{p}^{2}}(\|{\bm{M}}\|_{2}^{2}+\|{\bm{\Omega}}\|_{2}^{2})-\sum_{ij}\log\Omega_{ij}.$ (11)
@@ 3.4 On Improving the Sample Efficiency of BLoB
7:     Sample a mini-batch of data ${\mathcal{D}}_{t}\sim{\mathcal{D}}$.
@@
10:         Let $\widetilde{{\bm{E}}}_{k}\leftarrow\operatorname{BLoBFlipout}({\bm{E}}_{k})$$\triangleright$ Eqn. 12
11:         Let ${\bm{A}}_{k}={\bm{M}}+{\bm{G}}^{2}\circ\widetilde{{\bm{E}}}_{k}$.
@@
13:     Let ^$\widehat{{\mathcal{F}}}_{t}=-\tfrac{1}{K}\sum_{k=1}^{K}\log P({\mathcal{D}}_{t}|{\bm{A}}_{k},{\bm{B}})+\tfrac{1}{2\sigma_{p}^{2}}(\|{\bm{M}}\|_{2}^{2}+\|{\bm{G}}\|_{2}^{4})-2\sum_{ij}\log G_{ij}$. $\triangleright$ Eqn. 3.5 and 11.
14:     Calculate the gradient w.r.t. the parameters:  $\Delta_{{\bm{M}}}=\nicefrac{{\partial\widehat{{\mathcal{F}}}_{t}}}{{\partial{\bm{M}}}}$, $\Delta_{{\bm{G}}}=\nicefrac{{\partial\widehat{{\mathcal{F}}}_{t}}}{{\partial{\bm{G}}}}$, $\Delta_{{\bm{B}}}=\nicefrac{{\partial\widehat{{\mathcal{F}}}_{t}}}{{\partial{\bm{B}}}}$.
15:     Update the parameters:  ${\bm{M}}\leftarrow{\bm{M}}-\eta\Delta_{{\bm{M}}}$;  ${\bm{G}}\leftarrow{\bm{G}}-\eta\Delta_{{\bm{G}}}$;  ${\bm{B}}\leftarrow{\bm{B}}-\eta\Delta_{{\bm{B}}}$
16:end for
LoRA Flipout. Unlike the original approach, which applies rank-1 random flipping to the full weights, we apply flipout exclusively to the low-rank component ${\bm{A}}$. Specifically, suppose we have a mini-batch of input vectors ${\bm{H}}\in\mathbb{R}^{n\times b}$, where $b$ represents the batch size. We randomly sample two low-rank flipping matrices ${\bm{S}}\in\{-1,+1\}^{n\times b}$ and ${\bm{T}}\in\{-1,+1\}^{b\times r}$. Denoting as ${\bm{E}}$ the weight noise sampled for this mini-batch, the batched output ${\bm{Z}}$ after applying flipout is then
@@
where $\circ$ denotes the element-wise product (see Appendix A.3 for derivation). It is crucial that the independent noises added to the low-rank weight noise $\Delta{\bm{A}}\triangleq{\bm{E}}\circ{\bm{\Omega}}$ ensure sampling independence across examples within a mini-batch, thereby enhancing the sampling efficiency of the algorithm. This is done without violating the assumptions outlined in Theorem 3.1 and 3.2. Besides, due to the low-rank structure of our Bayesianization method, the computational overhead of employing flipout in BLoB is also minimal.
@@ 3.5 BLoB: Final Algorithm
We are now ready to present our full BLoB algorithm. Under the assumptions outlined in Theorem 3.1 and 3.2, optimizing the evidence lower bound on the full weight ${\bm{W}}$ can be efficiently done in the low-rank space, using the following final objective function:
@@
where ${\bm{\theta}}=\{{\bm{M}},{\bm{\Omega}}\}$ denotes the set of the parameters underlying the variational distribution of the low-rank matrix ${\bm{A}}$. The full algorithmic description of BLoB is shown in Algorithm 1.
@@ 4.1 Settings
Table 1: Performance of different methods applied to LoRA on Llama2-7B pre-trained weights. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5,000 training steps. We use $N$ to represent the number of samples during inference in BLoB. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
@@
LAP [108] 69.20$\pm$1.50 29.73$\pm$12.02 80.05$\pm$0.22 75.55$\pm$0.36 82.12$\pm$0.67 86.95$\pm$0.09
@@
LAP [108] 4.15$\pm$1.12 14.24$\pm$1.65 33.29$\pm$0.57 7.40$\pm$0.27 8.70$\pm$1.77 1.30$\pm$0.33
@@
LAP [108] 0.60$\pm$0.00 1.33$\pm$0.05 0.88$\pm$0.00 0.57$\pm$0.01 0.52$\pm$0.01 0.31$\pm$0.00
@@
Baselines and Implementation Details. We compare BLoB with state-of-the-art uncertainty estimation methods applied to the LoRA adapters of LLMs, including Maximum Likelihood Estimation (MLE[41], Maximum A Posteriori (MAP), Monte-Carlo Dropout (MCD[29], Bayes By Backprop (BBB[11], Deep Ensemble (ENS[51, 8, 96], and the latest Laplace-LoRA (LAP[108].
4.2 Results on In-distribution Datasets
@@
BLoB consistently achieves better or comparable performance across all datasets. With the number of samples during inference set to $N=10$, the same as MCD, BLoB provides the best uncertainty estimation performance, significantly reducing NLL and ECE, and greatly mitigating overconfidence while maintaining comparable or better ACC than MLE. Even with half the number of samples, $N=5$, BLoB still delivers performance comparable to that of $N=10$ and outperforms other baselines on most datasets. Surprisingly, using only the mean of the weight distribution, BLoB achieves significantly higher ACC than other baselines on smaller datasets (WG-S, ARC-C, ARC-E) without a substantial loss in uncertainty estimation quality. This observation further confirms that jointly learning the mean and covariance during fine-tuning can mutually improve their quality.

### v5

@@ Abstract
Large Language Models (LLMs) often suffer from overconfidence during inference, particularly when adapted to downstream domain-specific tasks with limited data. Previous work addresses this issue by employing approximate Bayesian estimation after the LLMs are trained, enabling them to quantify uncertainty. However, such post-training approaches’ performance is severely limited by the parameters learned during training. In this paper, we go beyond post-training Bayesianization and propose Bayesian Low-Rank Adaptation by Backpropagation (BLoB), an algorithm that continuously and jointly adjusts both the mean and covariance of LLM parameters throughout the whole fine-tuning process. Our empirical results verify the effectiveness of BLoB in terms of generalization and uncertainty estimation, when evaluated on both in-distribution and out-of-distribution data. Code is available at https://github.com/Wang-ML-Lab/bayesian-peft.
@@ 1 Introduction
Despite the recent advancements in Large Language Models (LLMs) [9, 106, 105, 65, 15, 4, 91, 92, 77, 12, 1, 71], addressing the challenges of reliability and responsibility remains imperative [41, 100, 99]. LLMs often produce overconfident responses detached from factual grounding, posing potential harm to users [3, 107, 43, 42, 87, 49, 7, 118, 111, 120, 32, 68, 115, 44]. Therefore, accurately estimating response confidence (or uncertainty) is crucial to preemptively intervene before harm occurs. Current research predominantly focuses on eliciting the internal capability of uncertainty estimation of LLMs. For example, studies suggest that verbalized uncertainty yields better-calibrated results compared to conditional probability [87, 44].
While effective, the aforementioned methods do not offer a universal solution for expressing LLM uncertainty across all scenarios, especially when adapted [113] to domain-specific corpora, human preferences, or downstream tasks [43]. Even a well-calibrated LLM may struggle to estimate uncertainty during fine-tuning due to catastrophic forgetting of general knowledge [84]. Moreover, when applied to limited-scale downstream tasks, excessively over-parameterized LLMs can rapidly overfit, leading to overconfidence. Thus, enabling accurate uncertainty estimation of LLMs is vital for their reliable and responsible deployment.
Bayesian methods emerge as a natural solution for learning uncertainty estimation abilities among their counterparts [88, 11, 101, 98, 29, 45, 50, 60, 97, 57, 102, 20, 110]. These methods model predictive uncertainty $P({\bm{y}}|{\bm{x}},{\mathcal{D}})$ by marginalizing the posterior parameter distribution $P({\bm{\theta}}|{\mathcal{D}})$ after observing the dataset ${\mathcal{D}}$:
@@ 2.2 Variational Bayesian Networks (VBNs)
${\mathcal{F}}({\mathcal{D}},{\bm{\theta}})\triangleq-\mathbb{E}_{q({\bm{W}}|{\bm{\theta}})}[\log P({\mathcal{D}}|{\bm{W}})]+\operatorname{KL}[q({\bm{W}}|{\bm{\theta}})\parallel P({\bm{W}})].$ (3)
@@
${\mathcal{F}}({\mathcal{D}},{\bm{\theta}})$ $\approx-\tfrac{1}{K}\sum_{k=1}^{K}\log P({\mathcal{D}}|{\bm{W}}_{k})+\tfrac{1}{K}\sum_{k=1}^{K}[\log q({\bm{W}}_{k}|{\bm{\theta}})-\log P({\bm{W}}_{k})],$ (4)
where ${\bm{W}}_{k}={\bm{\mu}}+\log(1+\exp({\bm{\rho}}))\odot{\bm{\epsilon}}_{k}$ is the $k$-th sample of the weights yielded by parameterization and ${\bm{\epsilon}}_{k}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$. In BBB, the authors assume the prior distribution $P({\bm{W}})=\pi{\mathcal{N}}({\bm{0}},{\bm{\sigma}}_{1}^{2})+(1-\pi){\mathcal{N}}({\bm{0}},{\bm{\sigma}}_{2}^{2})$ to be a mixture of Gaussians. Consequently, they optimize the second term based on weight sampling. In different scenarios, a simpler form of the prior, which allows for a closed-form solution, can also be considered. Although our proposed method is largely based on the existing framework of BBB, trivially combining BBB with LoRA does not yield satisfactory results. It is important to note that our specific designs are necessary to encourage the fast convergence of the variational distribution, which will be introduced later in Sec. 3.
@@ 3.1 Low-Rank Variational Approximate Posterior Distribution: LoRA Bayesianization
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}].$ (8)
Theorem 3.1 shows that our asymmetric LoRA Bayesianization is equivalent to using a Gaussian variational distribution for the full weight ${\bm{W}}$ (i.e., Eqn. 6), with a flexible covariance matrix (i.e., Eqn. 8), to approximate the posterior distribution of the full weight ${\bm{W}}$.
@@ 3.2 Low-Rank Prior Distribution
Theorem 3.2 shows that with a proper $\widetilde{{\bm{R}}}$, one can compute the KL divergence for the high-dimensional full weight $\operatorname{vec}({\bm{W}})$ simply by computing the KL divergence for ${\bm{A}}$, which is much lower-dimension, more parameter-efficient, more memory-efficient, and faster. Note that the Gaussian distributions we define for both the prior and the posterior are degenerate. However, they are valid for probabilistic inference [83], as (i) their probability density is well-defined, and (ii) their KL divergence is computable under the assumptions of Theorem 3.2. See Appendix A.1 for a detailed discussion.
@@
$\operatorname{KL}[q({\bm{A}}|{\bm{\theta}}=\{{\bm{M}},{\bm{\Omega}}\})\|P({\bm{A}})]$ $=\tfrac{1}{2\sigma_{p}^{2}}(\|{\bm{M}}\|_{2}^{2}+\|{\bm{\Omega}}\|_{2}^{2})-\sum_{ij}\log\Omega_{ij}.$ (11)
@@ 3.4 On Improving the Sample Efficiency of BLoB
7:     Sample a mini-batch of data ${\mathcal{D}}_{t}\sim{\mathcal{D}}$ containing $b$ samples.
@@
10:         Let $\{\widetilde{{\bm{E}}}_{kj}\}_{j=1}^{b}\leftarrow\operatorname{BLoBFlipout}({\bm{E}}_{k})$$\triangleright$ Eqn. 12
11:         Let $\{{\bm{A}}_{kj}={\bm{M}}+{\bm{G}}^{2}\circ\widetilde{{\bm{E}}}_{kj}\}_{j=1}^{b}$.
@@
13:     Let ^$\widehat{{\mathcal{F}}}_{t}=-\tfrac{1}{Kb}\sum_{k=1}^{K}\sum_{j=1}^{b}\log P({\mathcal{D}}_{t}|{\bm{A}}_{kj},{\bm{B}})+\tfrac{1}{2\sigma_{p}^{2}}(\|{\bm{M}}\|_{2}^{2}+\|{\bm{G}}\|_{2}^{4})-2\sum_{ij}\log G_{ij}$.
14$\triangleright$ Eqn. 3.5 and 11.
15:     Calculate the gradient w.r.t. the parameters:  $\Delta_{{\bm{M}}}=\nicefrac{{\partial\widehat{{\mathcal{F}}}_{t}}}{{\partial{\bm{M}}}}$, $\Delta_{{\bm{G}}}=\nicefrac{{\partial\widehat{{\mathcal{F}}}_{t}}}{{\partial{\bm{G}}}}$, $\Delta_{{\bm{B}}}=\nicefrac{{\partial\widehat{{\mathcal{F}}}_{t}}}{{\partial{\bm{B}}}}$.
16:     Update the parameters:  ${\bm{M}}\leftarrow{\bm{M}}-\eta\Delta_{{\bm{M}}}$;  ${\bm{G}}\leftarrow{\bm{G}}-\eta\Delta_{{\bm{G}}}$;  ${\bm{B}}\leftarrow{\bm{B}}-\eta\Delta_{{\bm{B}}}$
17:end for
LoRA Flipout. Unlike the original approach, which applies rank-1 random flipping to the full weights, we apply flipout exclusively to the low-rank component ${\bm{A}}$. Specifically, suppose we have a mini-batch of input vectors ${\bm{H}}\in\mathbb{R}^{n\times b}$, where $b$ represents the batch size. We randomly sample two low-rank flipping matrices ${\bm{S}}\in\{-1,+1\}^{n\times b}$ and ${\bm{T}}\in\{-1,+1\}^{b\times r}$. Denoting as ${\bm{E}}\in\mathbb{R}^{r\times n}$ the weight noise sampled for this mini-batch, the batched output ${\bm{Z}}$ after applying flipout is then
@@
It is crucial that the independent noises added to the low-rank weight noise $\Delta{\bm{A}}\triangleq{\bm{E}}\circ{\bm{\Omega}}$ ensure sampling independence across examples within a mini-batch, thereby enhancing the sampling efficiency of the algorithm. This is done without violating the assumptions outlined in Theorem 3.1 and 3.2. As illustrated in algorithm 1, we use $\widetilde{{\bm{E}}}_{kj}$ to represent the equivalent noise applied to parameter ${\bm{A}}$ for the $j$-th example in the $k$-th batch after BLoBFlipout. Due to the low-rank structure of our Bayesianization method, the computational overhead of employing flipout in BLoB is also minimal.
@@ 3.5 BLoB: Final Algorithm
We are now ready to present our full BLoB algorithm.
During training, under the assumptions outlined in Theorem 3.1 and 3.2, optimizing the evidence lower bound on the full weight ${\bm{W}}$ can be efficiently done in the low-rank space, using the following final objective function:
@@
where ${\bm{\theta}}=\{{\bm{M}},{\bm{\Omega}}\}$ denotes the set of the parameters underlying the variational distribution of the low-rank matrix ${\bm{A}}$. Additionally, to trade off between data fitting and posterior approximation, we employ a KL re-weighting scheme, which is detailed in Appendix B.1. The full algorithmic description of BLoB training is shown in Algorithm 1.
During inference, for an input ${\bm{x}}$, we approximate the expected output distribution $P({\bm{y}}|{\bm{x}})$ of BLoB by drawing $N$ samples from the variational distribution $q({\bm{W}}|{\bm{\theta}})$. Empirically, $N=10$ provides a good balance between estimation quality and computational efficiency:
$\mathbb{E}_{q({\bm{W}}|{\bm{\theta}})}[P({\bm{y}}|{\bm{x}},{\bm{W}})]$ $\approx\tfrac{1}{N}\sum_{n=1}^{N}P({\bm{y}}|{\bm{x}},{\bm{W}}_{n}),\quad{\bm{W}}_{n}\sim q({\bm{W}}|{\bm{\theta}}).$ (14)
@@ 4.1 Settings
Baselines and Implementation Details. We compare BLoB with state-of-the-art uncertainty estimation methods applied to the LoRA adapters of LLMs, including Monte-Carlo Dropout (MCD[29], Bayes By Backprop (BBB[11], Deep Ensemble (ENS[50, 8, 103], and the latest Laplace-LoRA (LAP[116]. We also report the performance of two standard PEFT baseline methods for reference: Maximum Likelihood Estimation (MLE[40] and Maximum A Posteriori (MAP).
Table 1: Performance of different methods applied to LoRA on Llama2-7B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5,000 gradient steps. We use $N$ to represent the number of samples during inference in BLoB. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
@@
LAP [116] 69.20$\pm$1.50 66.78$\pm$0.69111 LAP encounters training failures on the ARC-C dataset under a unified setting, where the same hyperparameter configuration is applied across all datasets. To achieve competitive performance with LAP on the ARC-C dataset, we deviate from this unified setting, allowing dataset-specific hyperparameters for LAP. Note that this introduces an unfair advantage for LAP. Specifically, for the LAP on ARC-C dataset, we set the dropout rate to 0.1, the learning rate to 5e-5, and applied early stopping at the 5,000-th iteration (out of 10,000). 80.05$\pm$0.22 75.55$\pm$0.36 82.12$\pm$0.67 86.95$\pm$0.09
@@
LAP [116] 4.15$\pm$1.12 16.25$\pm$2.61111 LAP encounters training failures on the ARC-C dataset under a unified setting, where the same hyperparameter configuration is applied across all datasets. To achieve competitive performance with LAP on the ARC-C dataset, we deviate from this unified setting, allowing dataset-specific hyperparameters for LAP. Note that this introduces an unfair advantage for LAP. Specifically, for the LAP on ARC-C dataset, we set the dropout rate to 0.1, the learning rate to 5e-5, and applied early stopping at the 5,000-th iteration (out of 10,000). 33.29$\pm$0.57 7.40$\pm$0.27 8.70$\pm$1.77 1.30$\pm$0.33
@@
LAP [116] 0.60$\pm$0.00 1.03$\pm$0.04111 LAP encounters training failures on the ARC-C dataset under a unified setting, where the same hyperparameter configuration is applied across all datasets. To achieve competitive performance with LAP on the ARC-C dataset, we deviate from this unified setting, allowing dataset-specific hyperparameters for LAP. Note that this introduces an unfair advantage for LAP. Specifically, for the LAP on ARC-C dataset, we set the dropout rate to 0.1, the learning rate to 5e-5, and applied early stopping at the 5,000-th iteration (out of 10,000). 0.88$\pm$0.00 0.57$\pm$0.01 0.52$\pm$0.01 0.31$\pm$0.00
@@
4.2 Results on In-distribution Datasets
BLoB consistently achieves better or comparable performance across all datasets. With the number of samples during inference set to $N=10$, the same as MCD, BLoB provides the best uncertainty estimation performance, significantly reducing NLL and ECE, and greatly mitigating overconfidence while maintaining comparable or better ACC than MLE. Even with half the number of samples, $N=5$, BLoB still delivers performance comparable to that of $N=10$ and outperforms other baselines on most datasets. By abandoning the modeling of the posterior distribution, prediction using the mean of the weight distribution, i.e., BLoB (N=0) sacrifices some degree of calibration in exchange for improved accuracy. Appendix C.5 presents the trade-off between accuracy and calibration, which is controlled by the standard deviation of the prior Gaussian distribution).